OpenStudy (anonymous):

the square root of 24

OpenStudy (anonymous):

4.89897948557

OpenStudy (anonymous):

simplifying expressions with complex numbers

jimthompson5910 (jim_thompson5910):

Hint: 24 = 4*6

jimthompson5910 (jim_thompson5910):

and \[\large \sqrt{x*y} = \sqrt{x}*\sqrt{y}\]